1 At a sales volume of 44,500 units, Choice Corporation's sales commissions (a cost that is...

1

At a sales volume of 44,500 units, Choice Corporation's sales commissions (a cost that is variable with respect to sales volume) total $640,800.

To the nearest whole cent, what should be the average sales commission per unit at a sales volume of 45,500 units? (Assume that this sales volume is within the relevant range.)

2.

Job 910 was recently completed. The following data have been recorded on its job cost sheet:

Direct materials $ 2,430
Direct labor-hours 70 labor-hours
Direct labor wage rate $ 20 per labor-hour
Machine-hours 134 machine-hours

The Corporation applies manufacturing overhead on the basis of machine-hours. The predetermined overhead rate is $21 per machine-hour. The total cost that would be recorded on the job cost sheet for Job 910 would be:

3.

Baka Corporation applies manufacturing overhead on the basis of direct labor-hours. At the beginning of the most recent year, the company based its predetermined overhead rate on total estimated overhead of $240,300 and 4,690 estimated direct labor-hours. Actual manufacturing overhead for the year amounted to $243,000 and actual direct labor-hours were 4,640.

The predetermined overhead rate for the year was closest to:

4.

Bayest Manufacturing Corporation uses a predetermined overhead rate based on direct labor-hours to apply manufacturing overhead to jobs. Last year, the Corporation worked 57,000 actual direct labor-hours and incurred $392,000 of actual manufacturing overhead cost. The Corporation had estimated that it would work 60,400 direct labor-hours during the year and incur $356,360 of manufacturing overhead cost. The Corporation's manufacturing overhead cost for the year was: (Round your intermediate calculations to 2 decimal places.
Answer #1

1) Average sales commissions per unit= $640800/44500= $14.40 per unit

2) Calculation of Total Cost

Direct materials $2430
Direct labor (70 labor hours*$20) 1400
Manufacturing overhead applied (134 machine hours*$21) 2814
Total manufacturing cost $6644

3) Predetermined overhead rate= Estimated total manufacturing cost/Estimated direct labor hours

= $240300/4690= $51.24 per direct labor hour

4) Predetermined overhead rate= Estimated total manufacturing cost/Estimated direct labor hours

= $356360/60400= $5.90 per direct labor hour

Manufacturing overhead applied= 57000*$5.90= $336300

Under or over applied= Actual manufacturing overhead-Manufacturing overhead applied

= $392000-336300= $55700 under applied

NOTE:- If you have any problem regarding the question please ask in the comment section.